OpenStudy (anonymous):

What are the odds that you will pick an ace from a standard deck of 52 cards?

OpenStudy (anonymous):

A.1 : 12 B.1 : 14 C.2 : 13 D.5 : 21

OpenStudy (kropot72):

The odds are the ratio of the probability of picking an ace to that of not picking an ace. There are 4 aces in the deck, so the probability of picking an ace is 4/52. Do you know the probability of not picking an ace?

OpenStudy (kropot72):

@WAyne5 Are you there?

OpenStudy (anonymous):

Yeah

OpenStudy (kropot72):

So what is the probability of not picking an ace?

OpenStudy (anonymous):

Idk but is the answer c because 52 divided 4 is 13 c is 2 : 13

OpenStudy (anonymous):

(4/52) / (48/52) = 1/12 so A

OpenStudy (anonymous):

are you sure it's C? odd = P(favor)/P(against) = 4/48 = 1/12

OpenStudy (anonymous):

Idk it was just a guess

OpenStudy (kropot72):

The probability of not picking an ace is: \[\frac{52-4}{52}=\frac{48}{52}\] So the odds are: \[\frac{4}{52}:\frac{48}{52}=1:12\]

OpenStudy (anonymous):

OK thanks for the help @sourwing and @kropot72

OpenStudy (kropot72):

You're welcome :)